OpenStudy (raeganj):

delores has colored cubes with the faces labeled 1-6. each face has a different number. if delores rolls the three cubes, what is the probability she will roll an even number on the white cube,a 3 on the red cube and a 5 on the blue cube?

OpenStudy (anonymous):

How many even numbers are there on the white cube?

OpenStudy (raeganj):

3

OpenStudy (anonymous):

Good 3 out of the 6 How many 3's are on the red cube?

OpenStudy (raeganj):

1

OpenStudy (anonymous):

Good 1 out of the 6 How many 5's are on the blue cube?

OpenStudy (anonymous):

Good so 3/6 * 1/6 * 1/6 = 3/216

OpenStudy (raeganj):

3/216? how?

OpenStudy (anonymous):

3 * 1 * 1 = 3 numerator 6 * 6 * 6 = 216 denominator

OpenStudy (raeganj):

it would have to be something that i can simplify all the answer choices on the hw are simplified

OpenStudy (anonymous):

OH 3 goes into 216 so it is 1/72

OpenStudy (raeganj):

oh, i feel dumb lol thanx hunn :)